OpenStudy (anonymous):

What is the slope of the line 4x - 8y -16 = 0 in the simplest form?

OpenStudy (anonymous):

the slope is 1/2

OpenStudy (anonymous):

how is this done?

OpenStudy (anonymous):

because: 4x-8y-16=0 4x-8y=16 -8y=16-4x y=-2+x/2 y=x\2 -2 once its its in y=mx +c form you take m and that is the slope therefore the slope is 1/2

OpenStudy (anonymous):

oh, that is totally great. Lifesaver